# Mapping the Podcast Ecosystem with the Structured Podcast Research Corpus

### v1

@@ Appendix A Transcription
After preliminary experiments with different model sizes, we settled on using the base version of Whisper, with 74M parameters, as a compromise between speed and accuracy. To enable highly parallelized throughput, we use a version adapted to run on CPUs.888https://github.com/ggerganov/whisper.cpp Altogether, we transcribed over 650K hours of audio, which required approximately roughly 220K CPU hours.
@@
Following a manual review of over 100 segments, we discover only one case in which text was hallucinated. In this case, the words in the transcript which did not occur in the audio appeared at the end of the transcript, with time stamps that extended beyond the length of the audio file, thus could easily be filtered out.
@@ Appendix B Diarization
Appendix C Speaker Role Labelling
Because hosts and guests are not identified in any systematic fashion, we developed a system to determine this automatically, where possible. In particular, we rely on the fact that hosts frequently identify themselves and their guests at the start of an episode. We begin by identifying candidate names by using spaCy to perform NER on the first 350 words of the transcript and in podcast and episode descriptions. We keep only Person entities, and only those that are two word phrase (rather than single names). We further note that episodes can have one or most hosts, and zero or more guests. Based on this setup, we develop and deploy a system to classify each extracted name as Host, Guest, or Neither.
@@ Appendix C Speaker Role Labelling
In order to train a model to perform this classification, we collect human judgements as to whether extracted names are Host, Guest, or Neither. We set this up as a classification task for humans, in which we show annotators a single target entity name, along with the necessary context (podcast description, episode description, and 150 words surrounding the name’s first appearance in the transcript). Using Prolific and the Potato annotation tool (Pei et al., 2022), we collect three human judgements on each of 2,000 entity names. The task consists of a sequence of categorical annotation questions, each of which shows occurrences of the same named-entity within podcast descriptions, episode descriptions, and episode transcripts.
@@
For our second round of annotation, we trained an initial host-guest model on our first-round annotations and took the 1,000 entities with the lowest-confidence predictions from a random sample of 15,000 unique entity predictions.
Our annotation pipeline included 3 annotations per entity. Chance corrected annotator agreement (using Krippenforff’s ) was 0.77 for the first round, and 0.53 for the second round, since the second round consisted of more difficult examples. To account for malicious annotation strategies (e.g. spamming), we used the MACE tool to predict the most probable annotation for each entity Hovy et al. (2013).
@@
To identify the names of hosts and guests for each episode, we apply our trained model to all the extracted entity names for that episode. Once again, the maximum-probability prediction was used for entities occurring multiple times in the same transcript. To evaluate the model, we only used entities from the first round of annotation in our held-out set, since they are a more representative sample. Our final model achieved a mean cross validation accuracy of .87 and a test accuracy of .88. Distributions of the number of inferred hosts and guests across categories are shown in Figures 5 and 6.
@@ Appendix D Data Merging and Dataset Release
We release our dataset in two formats corresponding to podcast and speaker-turn level information respectively. At the podcast level, we release full transcripts, podcast and episode level metadata, and inferred host and guest assignments. At the speaker-turn level, we provide aggregated prosodic features, inferred speaker roles and names, the transcribed speaker turn, and a unique key mapping the turn back to the podcast-level metadata. When identifying speaker turns for segments of words with overlapping text, we assign this text to the speaker who had already been speaking when the overlap occurred.
@@ Appendix E Descriptive Overview
Figure 7 summarizes statistics of our podcast dataset regarding length, hosting, and categories. We find that the median podcast length is 30 minutes, with approximately 80 percent of podcasts lasting one hour or less. Our data also show that podcast hosting is highly centralized, with anchor.fm (owned by Spotify) hosting nearly 6 times the number of podcasts as SoundCloud, the next largest host. Whereas podcasts began and continue to be a space for independent creators, this centralization is inline with prior work regarding the medium’s evolution as a source for corporate profit and mass-media communications Baldini (2015). Finally, podcast creators can assign their podcasts to a number of categories. We find that Religion is by far the most popular category (in terms of episodes), followed by Society, Education, Business, Sports, and News.
Appendix F Host-Guest Network
Constructing the network of guest co-appearances across podcasts requires identifying unique guests that appear on multiple shows. Unfortunately, based only on names, there is the potential for misidentifying all individuals with the same name as one person. To minimize the chance of this possibility, we filter out guests that have especially common names. We first estimate the probability of first and last names in our data separately, by splitting all identified named entities into two names (first and last), and computing the frequency of occurrence of each of these individually. We then model the overall probability of a two-word entity name as . This naively assumes that first and last name probabilities are independent, which they clearly are not, but is sufficient for our purposes. We then keep only those guests having below the median name probability across all guests.
Appendix G Calculating Modularity
For a set nodes partitioned into communities in a graph, modularity measures the difference between the number of edges that fall within their community minus the number that would be expected if the edges were reassigned randomly. More formally, modularity is given by equation (1), where indicates whether two nodes are connected based on adjacency matrix , is the (approximate) expected probability of an edge occurring between nodes and , and is an indicator variable for whether and are in the same community Newman (2006).
@@ Appendix H Topics
Appendix I Racial Justice Topics Across Categories

### v2

@@ Appendix A Transcription
Details
After preliminary experiments with different model sizes, we settled on using the base.en version of Whisper, with 74M parameters, as a compromise between speed and accuracy. To enable highly parallelized throughput, we use a version adapted to run on CPUs.101010https://github.com/ggerganov/whisper.cpp Altogether, we transcribed over 650K hours of audio, which required approximately roughly 220K CPU hours.
@@
Validation
To help quantify the performance of Whisper on podcasts, we compare the output on a small sample of episodes for which we can obtain professional transcripts shared by podcasts creators. Through web searches, we identify six shows that release transcripts for every episode on their websites, complete with a named speaker for each speaking turn. Of these shows, one (Welcome to Night Vale) typically consists of monologues, read by one speaker; another (This American Life) often involves many speakers over the course of one episode; and the remaining four are typically extended one-on-one interviews. From these shows, we download the transcripts for six episodes per show, starting from the beginning of May 2020, and taking each episode after that in turn.
Unfortunately, we cannot directly compare the output of Whisper to these transcripts, in part because the transcripts often exclude things like the intro and outro to the episode, and may be annotated with additional text throughout (including speaker names). Thus, we first process the transcript files to remove everything but the spoken words. We then manually edit the output of Whisper to remove everything before the first word and after the last word of the transcript.111111In none of these cases did we encounter hallucinated passages that were not in the audio files.
Given these preprocessed versions, for each episode, we tokenize both the transcript and the whisper output using spaCy, drop punctuation, and then use NLTK to align the two resulting lists of tokens based on edit distance. To compute word error rate (WER), we add up the number of insertions, deletions, and errors, and divide by the total number of tokens in the aligned list.
Our estimated WER values are show in Figure 4, with average WER per show shown by vertical bars. In most cases, the average WER is between 8-9%, which is better than the average WER reported in the original Whisper paper.121212The Whisper paper (Radford et al., 2022) reported an average WER of 12.8% for Whisper Large V2 across 13 datasets, with average WER per dataset varying between 3.9% and 36.4%. However, the error rates measured by this fully automated comparison actually overstate the true error rate by a considerable degree. The reason is that the professionally produced transcripts are both imperfect, and intentionally clean up the words spoken on the episodes, often removing disfluencies, like filler words (e.g., “um”, “like”, “you know”, etc.) and repetitions (e.g., “I mean, I mean”). Whisper does sometimes elide these as well, but is generally much closer to an exact transcription. The effect of this can be seen in the low WER for Welcome to Night Vale (average WER of 4.3%), which consists of professional readings of pre-written scripts (McGlynn, 2014), and thus contains far fewer of these disfluencies than natural speech.
To get a better estimate of the true transcription performance of Whisper, we manually annotate the alignment files for the first five minutes (from the start of the transcript) for each of eighteen episodes (the first three episodes from each show in our validation sample), and mark cases where Whisper was legitimately correct or incorrect, and with what kind of error. In doing these annotations, we are strict with respect to words being correct (so that “Amazons” is not equivalent to “Amazon’s”), but we do count equally plausible spellings of words and phrases as correct (except for proper names). This includes numbers (e.g., “11” vs. “eleven”), times (e.g,. “9:00pm” vs. “9 o’clock p.m”), money (e.g., “$279” vs. “two hundred seventy nine bucks”), alternative spellings (e.g., “advisor” vs. “adviser”). This also includes words that could plausibly be split (e.g., “best selling” vs. “bestselling”), and conjunctions and alternatives which cannot be easily decided from the audio (e.g,. “Today is” vs. “Today’s” or “going to” vs. “gonna”).
The results of this manual analysis are presented in Figure 5, and reveal that Whisper is actually performing much better than our automated analysis initially suggested. Based on our manual analysis of these eighteen episodes, we compute an average WER of 3.0%. However, it turns out that this even number is inflated somewhat by one particular episode of This American Life, which contains a high number of Spanish words that Whisper fails to translate, and which thus ends up with a WER of 12.6%. If we exclude this episode, then our manual analysis produces an estimated average WER for English of 2.4%, which is just under what the Whisper paper reported as performance on LibriSpeech Clean (2.7%).131313Note, however, that just as our automated analysis was too pessimistic, our manual analysis may be slightly too optimistic. The reason is that we do not penalize Whisper for omitting filler words or disfluencies that the professional transcripts also omit. These were quite rare, but including these would slightly increase the WER, depending on what we consider to be the “correct” transcription. Ultimately, mapping conversational speech to a clean sequence of tokens is a complex task, and may not be well-defined if only working at the level of tokens, especially in the case of overlapping speech.
Looking at the annotated alignments in more detail, we find that in cases where there is a disagreement about a word between Whisper and the professional transcript, we find that the Whisper is correct about half the time. In cases where Whisper produces the wrong word, this is most commonly due to proper names, which are legitimately ambiguous in terms of spelling, without additional world knowledge. In cases where the transcript is missing a word that is in the Whisper output, Whisper is correct that the word was spoken more than 80% of the time. When Whisper is missing a word that is in the transcript (which is the least common type of error), this is usually a legitimate error, where Whisper has failed to transcribe a word, which is often a filler word that was retained by the transcript.
Without further analysis, it is difficult to know how well these estimates will generalize to other shows. However, the fact that we observe error rates of around 4% or less on most episodes, across six shows with some diversity of speakers, suggests that transcription performance tends to be quite reliable overall.
@@ Appendix B Diarization
Validation
To validate the performance of pyannote on this data, we again make use of the professional transcripts described in Appendix A. Working with the raw diarization output (prior to filtering), we similarly manually edit these to remove everything before the first word of the transcript and everything after the last word of the transcript. For the preprocessed version of each of the professional transcript and the pyannote output, we again tokenize with spaCy, drop punctuation, and convert each to a sequence of tuples, where each tuple represents one token and the corresponding set of one or more speakers. For the professional transcripts, this is always a single named speaker per token. For the diarization output the speakers are unnamed (e.g., speaker1, speaker2), and the set might contain one speaker or more (in the case of inferred overlap).
To determine how well the two assignments of tokens to speakers align, we rely on the token alignments created during our transcription validation (see Appendix A). In particular, we use the alignment between the whisper output and the professional transcript to get the named speaker from the transcript for each token.141414Because of minor differences in tokenization which sometimes happen, we allow for matching sequences of tokens (rather than just individual tokens) as necessary. For each token in the diarization output, we thus have the (correct) named speaker, and a set of one or more unnamed speakers. In doing so, we produce co-occurrence statistics on how many times each named speaker from the professional transcript is matched to each unnamed speaker from diarization. Finally, we associated each named speaker with the one unnamed speaker with the highest number of associated tokens, who has not yet been assigned to a name speaker. We do this in order of most tokens spoken among the named speakers.151515Note that this means that in some cases, an unnamed speaker from diarization might not get matched to any named speaker from the transcript.
Using this alignment of one unnamed speaker from diarization for each named speaker from professional transcripts, we can thus determine the proportion of tokens that have been correctly assigned. We compute the diarization token error rate for each episode, where we count as an error any token assigned to multiple unnamed speakers, or assigned to the wrong unnamed speaker (i.e., anyone other than the one associated with the correct named speaker for that token).
The results of this are highly encouraging, as shown in Figure 6. With the exception of one show, the diarization error rate for all episodes is less than 5%, and for most episodes less than 2%. The one show for which diarization struggles more (This American Life), is much more challenging due to the higher number of speakers on each episode. In our sample of six episodes, all episodes featured at least 17 speakers, with an average of 23.5 per episode. Note that although Welcome to Night Vale typically features only a single speaker, getting less than 100% diarization accuracy is possible if pyannote incorrectly infers that there is more than one speaker present in the audio. If we exclude This American Life, the overall average diarization token error rate is 1.8%, or 2.1% for those episodes which feature one-on-one interviews.
Appendix C Speaker Role Labeling
Because hosts and guests are not identified in any systematic fashion, we developed a system to determine this automatically, where possible. In particular, we rely on the fact that hosts frequently identify themselves and their guests at the start of an episode. We begin by identifying candidate names by using spaCy to perform NER on the first 350 words of the transcript and in podcast and episode descriptions. We keep only Person entities, and only those that are two-word phrase (rather than single names). We further note that episodes can have one or most hosts, and zero or more guests. Based on this setup, we develop and deploy a system to classify each extracted name as Host, Guest, or Neither.
@@ Appendix C Speaker Role Labeling
In order to train a model to perform this classification, we collect human judgments as to whether extracted names are Host, Guest, or Neither. We set this up as a classification task for humans, in which we show annotators a single target entity name, along with the necessary context (podcast description, episode description, and 150 words surrounding the name’s first appearance in the transcript). Using Prolific and the Potato annotation tool (Pei et al., 2022), we collect three human judgments on each of 2,000 entity names. The task consists of a sequence of categorical annotation questions, each of which shows occurrences of the same named entity within podcast descriptions, episode descriptions, and episode transcripts.
@@
For our second round of annotation, we trained an initial host-guest model on our first round annotations and took the 1,000 entities with the lowest confidence predictions from a random sample of 15,000 unique entity predictions.
Our annotation pipeline included 3 annotations per entity. Chance corrected annotator agreement (using Krippendorff’s ) was 0.77 for the first round, and 0.53 for the second round, since the second round consisted of more difficult examples. To account for malicious annotation strategies (e.g. spamming), we used the MACE tool to predict the most probable annotation for each entity Hovy et al. (2013).
@@
To identify the names of hosts and guests for each episode, we apply our trained model to all the extracted entity names for that episode. Once again, the maximum probability prediction was used for entities occurring multiple times in the same transcript. To evaluate the model, we only used entities from the first round of annotation in our held out set, since they are a more representative sample. Our final model achieved a mean cross validation accuracy of .87 and a test accuracy of .88. Distributions of the number of inferred hosts and guests across categories are shown in Figures 8 and 9.
@@ Appendix D Data Merging and Dataset Release
We release our dataset in two formats corresponding to podcast and speaker-turn level information respectively. At the podcast level, we release full transcripts, podcast and episode level metadata, and inferred host and guest assignments. At the speaker-turn level, we provide aggregated prosodic features, inferred speaker roles and names, the transcribed speaker turn, and a unique key mapping the turn back to the podcast-level metadata. When identifying speaker turns for segments of words with overlapping speaker labels, we assign this text to the speaker who had already been speaking when the overlap occurred.
@@ Appendix E Descriptive Overview
Figure 10 summarizes statistics of our podcast dataset regarding length, hosting, and categories. We find a median podcast length of 30 minutes, with approximately 80 percent of podcasts lasting one hour or less. Our data also show that podcast hosting is highly centralized, with anchor.fm (owned by Spotify) hosting nearly 6 times the number of podcasts as SoundCloud, the next largest host. This centralization is in line with prior work regarding the medium’s evolution as a source for corporate profit and mass media communications Bonini (2015). Finally, podcast creators can assign their podcasts to a number of categories. We find that Religion is by far the largest category, followed by Society, Education, Business, Sports, and News.
Appendix F Speech Characteristics
Although it is not the focus of our work, we provide one example of the potential uses of extracted audio information. Using the fundamental frequency (F0) values from OpenSMILE, we compute the average pitch per category of podcasts in our sample.
As expected, podcasts in the Kids category (aimed at children) have the highest average pitch. By contrast, we find the lowest average pitch in Science, Leisure, and Technology categories, a pattern which may reflect a relative lack of gender diversity (see Figure 11). Future analyses could look at variation within these categories, and other speech properties such as dialect and speaking rate, as well as additional vocal characteristics measured by OpenSMILE.
Appendix G Host-Guest Network
Constructing the network of guest co-appearances across podcasts requires identifying unique guests that appear on multiple shows. Unfortunately, based only on names, there is the potential for misidentifying all individuals with the same name as one person. To minimize the chance of this possibility, we filter out guests with common names. We first estimate the probability of first and last names in our data separately, by splitting all identified named entities into two names (first and last), and computing the frequency of occurrence of each of these individually. We then model the overall probability of a two-word entity name as . This naively assumes that first and last name probabilities are independent, which they clearly are not, but is sufficient for our purposes. We then keep only those guests having below the median name probability across all guests.
Appendix H Calculating Modularity
For a set of nodes partitioned into communities in a graph, modularity measures the difference between the number of edges that fall within their community minus the number that would be expected if the edges were reassigned randomly. More formally, modularity is given by equation (1), where indicates whether two nodes are connected based on adjacency matrix , is the (approximate) expected probability of an edge occurring between nodes and , and is an indicator variable for whether and are in the same community Newman (2006).
@@ Appendix I Topics
Appendix J Racial Justice Topics Across Categories